OpenStudy (anonymous):

Write and solve the proportion: 3 is to 12 as w is to 8. a) w=2 b) w=4 c) w=24 d) w=32

ganeshie8 (ganeshie8):

3:12 = w:8

ganeshie8 (ganeshie8):

right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you can write the same as : \(\huge \frac{3}{12} = \frac{w}{8}\)

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

now you need to solve for okay...

OpenStudy (anonymous):

its 2 then ?

ganeshie8 (ganeshie8):

solve for w

ganeshie8 (ganeshie8):

whats 2 ? could u be explicit... :)

OpenStudy (anonymous):

the answer is 2 right ? lol

ganeshie8 (ganeshie8):

nope... lets go step by step ok ?

OpenStudy (anonymous):

okay

ganeshie8 (ganeshie8):

hey you're right!!! answer is 2, w = 8*3/12 = 24/12 = 2 gw =))

OpenStudy (anonymous):

(: thank you @ganeshie8

ganeshie8 (ganeshie8):

np...... yw :))